OpenStudy (anonymous):

Quick medal! What is the x-coordinate of the solution to the system of equations?

OpenStudy (alexandervonhumboldt2):

which one?

OpenStudy (anonymous):

hold on

OpenStudy (alexandervonhumboldt2):

solution is where lines intersect

OpenStudy (anonymous):

x=y+16 y+ 4x =9

OpenStudy (alexandervonhumboldt2):

substitue x into equation y+4x=9

OpenStudy (anonymous):

so like y+4=9?

OpenStudy (alexandervonhumboldt2):

woorps i made a msiaike

OpenStudy (alexandervonhumboldt2):

y+4*(y+16)=9

OpenStudy (alexandervonhumboldt2):

here

OpenStudy (alexandervonhumboldt2):

y+4y+64=9

OpenStudy (alexandervonhumboldt2):

y=-11

OpenStudy (alexandervonhumboldt2):

now substitue y into one of the equartion and solve for x

OpenStudy (alexandervonhumboldt2):

the value of x is the asnwer

OpenStudy (anonymous):

What does x=

OpenStudy (alexandervonhumboldt2):

x=-11+16=5

OpenStudy (anonymous):

Thank you

OpenStudy (alexandervonhumboldt2):

np